# Isometries of spacetimes without observer horizons

### v1

@@ 1. Introduction
In addition to causality (i.e. antisymmetry of $\leq$), we assume the “no observer horizons” condition (NOH), which prevents the lightcones from becoming excessively narrow towards infinity. Equivalently, it ensures that every observer is able to receive signals from every point in spacetime (after waiting long enough). This condition has been extensively studied in the context of the causal boundary and Bartnik’s splitting conjecture [13]. As pointed out by Penrose [31], the combination of causality and the NOH implies another standard causality assumption, namely global hyperbolicity with compact Cauchy surfaces. It has been shown by various authors that spacetimes with certain kinds of “time-translation like” symmetry must satisfy the NOH [8, 11, 16]. Here, we turn the situation around, and study how the NOH restricts the possible symmetries of the spacetime.
@@
Let $(M,g)$ be a causal spacetime satisfying the no observer horizons condition. Then the group $\operatorname{Isom}^{\uparrow}(M,g)$ of time orientation preserving isometries acts properly on $M$.
@@
Properness of the action (which we recall later on) means essentially the action has a non-chaotic dynamics, so that all orbits are closed, and even more, the orbit space is Hausdorff. Corollary 1.3 shows that, when the NOH is satisfied, the splitting of the spacetime into a product of time and space via a Cauchy temporal function (due to Geroch [17] and Bernal and Sánchez [5]) induces a splitting of the isometry group. The NOH is essential here; a counterexample is given by de Sitter spacetime, which has compact Cauchy surfaces, but does not satisfy any of our conclusions. See also Müller [27] for prior work on temporal functions invariant under a Lie group action, where the spacetime is only assumed globally hyperbolic, but the Lie group is assumed compact.
@@ 2.1. Lie groups
Now let $G$ be a Lie group. Its Lie algebra $\mathfrak{g}$ is the vector space of smooth left-invariant vector fields on $G$, equipped with the usual Lie bracket [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ]. There is a canonical isomorphism of vector spaces $\mathfrak{g}\to T_{e}G$ between the Lie algebra and the tangent space at the identity $e\in G$. The exponential map $\exp\colon\mathfrak{g}\to G$ sends a vector $X\in\mathfrak{g}$ to $\exp(X):=\Phi(1)\in G$, where $\Phi\colon\mathbb{R}\to G$ is the unique one-parameter subgroup with $\Phi^{\prime}(0)=X$. A word of caution is appropriate here: We will also use $\exp$ to denote the exponential map on a semi-Riemannian manifold (defined through geodesics); it should be clear from context which one is meant.
@@
$[X,Y]=d\Psi(X)(Y)\text{ for }X\in\mathfrak{l},\ Y\in\mathfrak{n}.$
@@ 2.2. Group actions
The rest of the proof consists in arguing that an action that preserves a Riemannian metric and has proper orbit maps is proper. Suppose that $F_{i}\in O(M)$, $\phi_{i}\in\operatorname{Aut}(\Phi)$ are sequences such that $F_{i}\to F$ and $\phi_{i}(F_{i})\to\tilde{F}$ converge in $O(M)$. We need to show that $\phi_{i}$ has a convergent subsequence in $\operatorname{Aut}(\Phi)$ (see Proposition 2.1). Choose $r>0$ be small enough so that the $h$-balls $B_{r}(F)$ and $B_{4r}(\tilde{F})$ are compact. Let $i_{0}>0$ be large enough such that $d(F_{i},F)<r$ and $d(\phi_{i}(F_{i}),\tilde{F})<r$ for all $i\geq i_{0}$, where $d$ is the distance induced by $h$. Then
$\phi_{i}(F)\in\phi_{i}(B_{r}(F))\subseteq\phi_{i}(B_{2r}(F_{i}))=B_{2r}(\phi_{i}(F_{i}))\subseteq B_{4r}(\tilde{F}).$
Hence the sequence $\phi_{i}$ is contained in $\Theta_{F}^{-1}(B_{4r}(\tilde{F}))$, and the latter set is compact by properness of the orbit map $\Theta_{F}$. It follows that $\phi_{i}$ has a convergent subsequence, and hence the action is proper. ∎
@@ 2.3. Lorentzian geometry
It is equivalent to require the same conditions only on timelike curves. Moreover, the NFOH (NPOH) is equivalent to the future (past) causal boundary consisting of a single point. We formulate most results for the NFOH, but there is always an analogous version for the NPOH. In some cases, it is necessary to assume the full NOH. We first prove that the NFOH upgrades one of the lowest levels on the causal ladder, causality, to the highest level, global hyperbolicity (see also [31, p. 624] and [9, Sec. 9]).
@@
$(i)\implies(ii)$. Let $\tau$ be a Cauchy temporal function, $s\in\mathbb{R}$, and let $\Sigma$ denote $\tau^{-1}(s)$. By Theorem 2.5, $M$ splits into ×$\mathbb{R}\times\Sigma$. By (i), since $\lambda\mapsto(\lambda,x)$ is an inextendible causal curve for every $x\in\Sigma$, the collection $I^{-}((\lambda,x))\cap\Sigma$ for all $\lambda\in\mathbb{R}$ forms an open cover of $\Sigma$. By compactness of $\Sigma$, there is a finite subcover, and hence a value $T_{x}$ such that $\Sigma\subset I^{-}((T_{x},x))$. Property (ii) then follows if we can choose $T_{x}$ independent of $x$.
To see the latter, let $x\in\Sigma$ be fixed, and notice that if $y\in\Sigma\cap I^{-}((T_{x},x)))$, then there exist neighborhoods $U$ of $y$ and $V$ of $(T_{x},x)$ such that every point in $U$ is in the chronological past of every point in $V$. By compactness, we may cover $\Sigma$ by finitely many such $U$, and take the intersection $\tilde{V}$ of the corresponding $V$’s. Let $p\colon\mathbb{R}\times\Sigma\to\Sigma$ denote the canonical projection. Then, for all $z\in p(\tilde{V})$, we have that $\Sigma\subset I^{-}((T_{x},z))$. Repeating this construction for every $x\in\Sigma$, we obtain an open cover of $\Sigma$ by sets of the form $p(\tilde{V}_{i})$, and corresponding values $T_{i}$ such that $\Sigma\subset I^{-}((T_{i},z))$ for all $z\in p(\tilde{V}_{i})$. By compactness, there is a finite subcover $p(\tilde{V}_{j})$ for $j=1,...,J$. Setting $T=\max_{j}\{T_{j}\}$ we conclude that $\Sigma$ and $\tau^{-1}(T)$ are totally timelike connected. ∎
@@ 3.1. Proofs of the main results
Suppose that we have a converging sequence of frames $F_{i}$ at $p_{i}\to p$ such that $D\phi_{i}(F_{i})$ has converging basepoints but becomes degenerate in the limit. This can only happen if there is a sequence of spacelike vectors $v_{i}$ at $p_{i}$ that converges to a spacelike $v\neq 0$, such that $w_{i}:=c_{i}D\phi_{i}(v_{i})$ converges to a future-directed null vector $w$ at $q:=\lim_{i\to\infty}\phi_{i}(p_{i})$. Here $c_{i}$ is a sequence of positive numbers, which necessarily tends to 00, chosen so that $w_{i}$ converges (for instance, the $c_{i}$’s can be normalization constants with respect to an auxiliary Riemannian metric). This situation is depicted in Figure 1.
@@
Corollary 1.2 now arises as a special case of the following theorem. Note that here we do require the full NOH (and not just the NFOH).
@@
Let $(M,g)$ be a causal spacetime satisfying the NOH. Suppose that $G$ is a Lie group acting properly on $M$ by time-orientation preserving conformal transformations. Then
@@
(i) By Proposition 2.7 and Theorem 2.5, there is a Cauchy temporal function $\hat{\tau}\colon M\to\mathbb{R}$, with level sets denoted by $\Sigma_{t}:=\hat{\tau}^{-1}(t)$. Furthermore, let $f\colon\mathbb{R}\to\mathbb{R}$ a smooth function such that $f(x)=0$ for $x\leq-1$, $f^{\prime}(x)>0$ for $x\in(-1,1)$, and $f(x)=1$ for $x\geq 1$. Then $\tau_{0}=f\circ\hat{\tau}$ is a smooth function on $M$ satisfying
  • $\tau_{0}(p)=0$ for all $p\in J^{-}(\Sigma_{-1})$,
  • $d\tau_{0}$ is future-directed timelike on $I^{+}(\Sigma_{-1})\cap I^{-}(\Sigma_{1})$, and
  • $\tau_{0}(p)=1$ for all $p\in J^{+}(\Sigma_{1})$.
We start by defining what will be the gradient of our invariant temporal function by averaging with respect to the Haar measure $\mu$ on $G$,
[…]
By definition of $\tau_{0}$, it is equivalent to integrate over
$\hat{G}_{p}:=\{\phi\in G\mid\phi^{-1}(p)\in I^{+}(\Sigma_{-1})\cap I^{-}(\Sigma_{1})\}.$
Because $I^{+}(\Sigma_{-1})\cap I^{-}(\Sigma_{1})$ is compact and the action of $G$ is proper, we have that for every $p\in M$, $\hat{G}_{p}$ is a compact subset of $G$. Thus the integral is finite. Moreover,
$\Lambda_{p}:=\{\phi^{*}d\tau_{0}\mid\phi\in\hat{G}_{p}\}\subset T^{*}_{p}M$
@@
for every curve $c:[0,1]\to M$. The right hand side vanishes for closed curves, because for every $\phi$, $\phi^{*}d\tau_{0}$ is exact. It follows that $d\tau$ is also exact, and hence indeed it is the differential of a temporal function $\tau$.
It remains to show that $\tau$ is a Cauchy temporal function (at least, up to postcomposition with a real function). By the NOH and Lemma 2.10, the non-empty $\tau$-level sets are Cauchy surfaces. We want $\tau$ to attain all values in $\mathbb{R}$. Note that in any case, $\tau(\mathbb{R})$ is a connected interval, because $\tau$ is continuous and $\tau(\mathbb{R})=\tau(\gamma(\mathbb{R}))$ for $\gamma\colon\mathbb{R}\to M$ any inextendible causal curve. By statement (ii), elements of $G$ act by translations (proven below, where Cauchyness of $\tau$ is only needed to show properness, so the argument is not circular). If all elements act by the trivial translation, we may postcompose $\tau$ with a smooth, surjective, strictly increasing function $f\colon\tau(\mathbb{R})\to\mathbb{R}$, so that $f\circ\tau$ is an invariant temporal function that attains all values in $\mathbb{R}$. If, on the other hand, there is an element $\phi\in G$ that acts by a non-trivial translation, then $\tau\circ\phi$, and hence $\tau$ itself, must be unbounded.
@@
Let $(M,g)$ be a causal spacetime satisfying the NOH, and let $\tau\colon M\to\mathbb{R}$ be as in Theorem 3.2. Then
@@
has Lie algebra equal to $\mathfrak{g}$. Because $\tilde{G}$ is simply connected, by [20, Prop. 1.100], $\tilde{G}$ must be the universal covering group of $G$. Hence $G=\tilde{G}/\Gamma$ for $\Gamma\subset\tilde{G}$ a discrete central subgroup [20, Prop. 1.101]. Note that either $L\subset Z(\tilde{G})$ (if $\psi$ and hence $\Phi$ are trivial) or $L\cap Z(\tilde{G})=\{e\}$ (else). In the first case, we have nothing to prove, so assume that we are in the second one, implying $\Gamma\cap L=\{e\}$. Let $e_{S}\in S$, $e_{L}\in L$ denote the identity element, define
@@ 3.2. Examples
Claim. For every $x\in\Sigma$, the vertical line ×$\mathbb{R}\times\{x\}\subset M$ is preserved by $\phi$.
@@
In the above proof, compactness of $\Sigma$ is essential. Indeed, Minkowski spacetime $\mathbb{R}^{1,n}$ is also a product spacetime, but its isometry group, the Poincaré group, includes boosts (which mix space and time), and hence is larger than ×$\mathbb{R}\times\mathsf{O}(n)$. Note also that product spacetimes with compact slices always satisfy the NOH. An example where the NOH is violated and the conclusion of Corollary 1.3 fails, is given by de Sitter spacetime $\mathsf{dS}^{4}$. On the one hand, $\mathsf{dS}^{4}$ can be represented as the warped product $\mathsf{dS}^{4}=\mathbb{R}^{1,0}\times_{\cosh^{2}}S^{3}$, implying that is is globally hyperbolic with compact Cauchy surfaces [4, pp. 183–184]. On the other hand, $\mathsf{dS}^{4}$ can also be represented as the pseudosphere in Minkowski spacetime $\mathbb{R}^{1,4}$. It follows that its isometry group is the Lorentz group $\mathsf{O}^{\uparrow}(1,4)$ [29, Prop. 9.8], which, again because it contains boosts, cannot be written in the form $\mathbb{R}\ltimes N$ with $N$ compact. While the obvious examples of spacetimes with isometry group $\mathbb{R}\ltimes N$ are the ones that admit a complete timelike Killing vector field (such as $\partial_{t}$ in the case of a product spacetime), the following example shows that they are not the only ones.
@@
Consider on $\mathbb{R}^{2}$ the Lorentzian metric $g:=f(y)dx^{2}-dy^{2}$ for some function $f\colon\mathbb{R}\to[0,1)$ with $f(y)=f(y+1)$. Let $M$ be the quotient of $\mathbb{R}^{2}$ by the map $(x,y)\mapsto(x+1,y+1)$. The metric $g$ descends to the quotient, which we can model as ×$\mathbb{R}\times[0,1]$ with the identifications $(0,y)\sim(1,y+1)$. Hence $M$ is homeomorphic to a cylinder. Because $f<1$, there are no closed causal curves, and in fact, $(M,g)$ is globally hyperbolic with $\{(s,s)\mid s\in[0,1]\}/\sim$ being a compact Cauchy surface. Moreover, $(M,g)$ satisfies the NOH. The coordinate vector field $\partial_{x}$ is complete, spacelike and Killing, with integral curves escaping to infinity. For generic $f$, it is the only Killing vector field; hence $\operatorname{Isom}^{\uparrow,0}(M,g)\cong\mathbb{R}$ even though $(M,g)$ does not admit a timelike Killing vector field.
@@
Finally, we argue that $\operatorname{Isom}^{\uparrow}(M,g)=L\ltimes N$, where $L$ is generated by $\phi$, is a semi-direct product, and is not isomorphic to ×$\mathbb{Z}\times N$. Clearly, $\phi$ does not commute with all elements in $N$. Concretely, if $\chi=(\chi_{1},\chi_{2})\in\mathsf{O}(2)\times\mathsf{O}(2)$, then
@@
Let $(M,g)$ be a causal spacetime satisfying the NOH, and let $\tau\colon M\to\mathbb{R}$ be a Cauchy temporal function. Suppose $G$ is a group acting by isometries on $(M,g)$ such that, for every $t\in\mathbb{R}$ and one (equivalently, all) points $p\in\Sigma_{t}$, the orbit $G(p):=\{\phi(p)\mid\phi\in G\}$ equals $\Sigma_{t}$. Then
@@
Let $k\geq 2$ and let $\beta\colon\mathbb{R}\to\operatorname{GL}(k,\mathbb{R}),\ t\mapsto\beta_{t}$ be a one-parameter subgroup such that $\beta_{1}(\mathbb{Z}^{k})=\mathbb{Z}^{k}$ but $\beta_{1}\neq\operatorname{Id}$. On ×$\mathbb{R}\times\mathbb{R}^{k}$, consider the equivalence relation

### v2

@@ 1. Introduction
In addition to causality (i.e. antisymmetry of $\leq$), we assume the “no future observer horizons” condition (NFOH), which prevents the lightcones from becoming excessively narrow towards infinity. Equivalently, it ensures that every observer is able to receive signals from every point in spacetime (after waiting long enough). The precise definition is that $I^{+}(\gamma)=M$ for every future-inextendible causal curve $\gamma$. This condition has been extensively studied in the context of the causal boundary and Bartnik’s splitting conjecture [13]. As pointed out by Penrose [31], the combination of causality and the NFOH implies another standard causality assumption, namely global hyperbolicity with compact Cauchy surfaces. It has been shown by various authors that spacetimes with certain kinds of “time-translation like” symmetry must satisfy the NFOH [8, 11, 16]. Here, we turn the situation around, and study how the NFOH restricts the possible symmetries of the spacetime.
@@
Let $(M,g)$ be a causal spacetime satisfying the no future observer horizons condition. Then the group $\operatorname{Isom}^{\uparrow}(M,g)$ of time orientation preserving isometries acts properly on $M$.
@@
Properness of the action (which we recall later on) means essentially the action has a non-chaotic dynamics, so that all orbits are closed, and even more, the orbit space is Hausdorff. Corollary 1.3 shows that, when the NFOH is satisfied, the splitting of the spacetime into a product of time and space via a Cauchy temporal function (due to Geroch [17] and Bernal and Sánchez [5]) induces a splitting of the isometry group. The NFOH is essential here; a counterexample is given by de Sitter spacetime, which has compact Cauchy surfaces, but does not satisfy any of our conclusions. See also Müller [27] for prior work on temporal functions invariant under a Lie group action, where the spacetime is only assumed globally hyperbolic, but the Lie group is assumed compact.
@@ 2.1. Lie groups
Now let $G$ be a Lie group. Its Lie algebra $\mathfrak{g}$ is the vector space of smooth left-invariant vector fields on $G$, equipped with the usual Lie bracket $\cdot$$\cdot$]. There is a canonical isomorphism of vector spaces $\mathfrak{g}\to T_{e}G$ between the Lie algebra and the tangent space at the identity $e\in G$. The exponential map $\exp\colon\mathfrak{g}\to G$ sends a vector $X\in\mathfrak{g}$ to $\exp(X):=\Phi(1)\in G$, where $\Phi\colon\mathbb{R}\to G$ is the unique one-parameter subgroup with $\Phi^{\prime}(0)=X$. A word of caution is appropriate here: We will also use $\exp$ to denote the exponential map on a semi-Riemannian manifold (defined through geodesics); it should be clear from context which one is meant.
@@
$X,Y]=d\Psi(X)(Y)\text{ for }X\in\mathfrak{l},\ Y\in\mathfrak{n}.$
@@ 2.2. Group actions
[…]
Hence the sequence $\phi_{i}$ is contained in $)\Theta_{F}^{-1}(B_{4r}(\tilde{F}$)), and the latter set is compact by properness of the orbit map $\Theta_{F}$. It follows that $\phi_{i}$ has a convergent subsequence, and hence the action is proper. ∎
@@ 2.3. Lorentzian geometry
It is equivalent to require the same conditions only on timelike curves. Moreover, the NFOH (NPOH) is equivalent to the future (past) causal boundary consisting of a single point. We formulate most results for the NFOH, but there is always an analogous version for the NPOH. We first prove that the NFOH upgrades one of the lowest levels on the causal ladder, causality, to the highest level, global hyperbolicity (see also [31, p. 624] and [9, Sec. 9]).
@@
$(i)\implies(ii)$. Let $\tau$ be a Cauchy temporal function, $s\in\mathbb{R}$, and let $\Sigma$ denote $\tau^{-1}(s)$. By Theorem 2.5, $M$ splits into $\mathbb{R}\times\Sigma$. By (i), since $\lambda\mapsto(\lambda,x)$ is an inextendible causal curve for every $x\in\Sigma$, the collection $I^{-}((\lambda,x))\cap\Sigma$ for all $\lambda\in\mathbb{R}$ forms an open cover of $\Sigma$. By compactness of $\Sigma$, there is a finite subcover, and hence a value $T_{x}$ such that $\Sigma\subset I^{-}((T_{x},x))$. Property (ii) then follows if we can choose $T_{x}$ independent of $x$.
[…]
@@ 3.1. Proofs of the main results
Suppose that we have a converging sequence of frames $F_{i}$ at $p_{i}\to p$ such that $D\phi_{i}(F_{i})$ has converging basepoints but becomes degenerate in the limit. This can only happen if there is a sequence of spacelike vectors $v_{i}$ at $p_{i}$ that converges to a spacelike $v\neq 0$, such that $w_{i}:=c_{i}D\phi_{i}(v_{i})$ converges to a future-directed null vector $w$ at $q:=\lim_{i\to\infty}\phi_{i}(p_{i})$. Here $c_{i}$ is a sequence of positive numbers, which necessarily tends to 0, chosen so that $w_{i}$ converges (for instance, the $c_{i}$’s can be normalization constants with respect to an auxiliary Riemannian metric). This situation is depicted in Figure 1.
@@
Corollary 1.2 now arises as a special case of the following theorem.
@@
Let $(M,g)$ be a causal spacetime satisfying the NFOH. Suppose that $G$ is a Lie group acting properly on $M$ by time-orientation preserving conformal transformations. Then
@@
(i) By Proposition 2.7 and Theorem 2.5, there is a Cauchy temporal function $\hat{\tau}\colon M\to\mathbb{R}$. Let $\mu$ be a left-invariant Haar measure on $G$. If $G$ is compact, then $\mu(G)$ is finite, and hence
(4) $d\tau:=\int_{G}\phi^{*}d\tau_{0}\ d\mu(\phi).$
is a well-defined one-form on $M$. Moreover,
$\Lambda_{p}:=\{\phi^{*}d_{p}\hat{\tau}\mid\phi\in G\}\subset T^{*}_{p}M$
@@
for every curve $c:[0,1]\to M$. The right hand side vanishes for closed curves, because for every $\phi$, $\phi^{*}d\hat{\tau}$ is exact. It follows that $d\tau$ is also exact, and hence indeed it is the differential of a temporal function $\tau$.
It remains to deal with the case of non-compact $G$, the problem being that the integral (4) is then infinite. To avoid this, we replace $\hat{\tau}$ by a suitable function $\tau_{0}$ with compact support $K\subset M$. Then,
(5) $d\tau:=\int_{G}\phi^{*}d\tau_{0}\ d\mu(\phi)=\int_{G_{p}}\phi^{*}d\tau_{0}\ d\mu(\phi),$
where
$G_{p}=\{\phi\in G\mid\phi^{-1}(p)\in K\}$
is compact in $G$ (by propreness of the action), and hence the integral (5) is finite.
To define $\tau_{0}$, let $a<b$ and let $f\colon\mathbb{R}\to\mathbb{R}$ be a smooth function such that $f(x)=0$ for $x\leq a$, $f^{\prime}(x)>0$ for $x\in(a,b)$, and $f(x)=1$ for $x\geq b$. Then $\tau_{0}=f\circ\hat{\tau}$ is a smooth function on $M$ satisfying
  • $\tau_{0}(p)=0$ for all $p\in J^{-}(\Sigma_{a})$,
  • $d\tau_{0}$ is future-directed timelike on $K:=I^{+}(\Sigma_{a})\cap I^{-}(\Sigma_{b})$, and
  • $\tau_{0}(p)=1$ for all $p\in J^{+}(\Sigma_{b})$.
In order for (5) to be everywhere timelike, we need to ensure that every $G$-orbit has a representative in $K$ (because if there is $p$ such that the orbit of $p$ is disjoint from $K$, then $d_{p}\tau=0$, which we have to avoid).
By Lemma 2.9, there are real numbers $a<b<c<d<e$ such that the corresponding $\hat{\tau}$ level sets $\Sigma_{a}$, $\Sigma_{b}$, $\Sigma_{c}$, $\Sigma_{d}$, and $\Sigma_{e}$ are pairwise totally timelike connected. Consider the compact set $A:=\hat{\tau}^{-1}((a,e))=I^{+}(\Sigma_{a})\cap I^{-}(\Sigma_{e})$ and a point $p\in\Sigma_{c}\subset A$. By propreness of the action, if $G$ is non-compact, then there must be $\phi\in G$ such that $\phi(p)\not\in A$. Since $\Sigma_{c}$ is a Cauchy surface, so is $S_{1}:=\phi(\Sigma_{c})$, and it follows by achronality that $S_{1}\cap\hat{\tau}^{-1}((b,d))=\emptyset$. Therefore $\Sigma_{c}$ and $S_{1}$ are totally timelike connected (or $S_{1}$ and $\Sigma_{c}$, but then we may replace $\phi$ by $\phi^{-1}$, so there is no loss in generality). Moreover, we recursively construct a family of Cauchy surfaces $S_{0}:=\Sigma_{c}$, $S_{i}:=\phi^{i}(S_{0})$ for all $i\in\mathbb{Z}$ with $S_{i+1}\subset I^{+}(S_{i})$. We claim that every $G$-orbit contains a representative in the compact set
(6) $K_{i}:=J^{+}(S_{i})\cap J^{-}(S_{i+1}),$
for, say, $i=0$ (compactness follows from global hyperbolicity, given that $S_{i}$ is compact).
To see that every $G$-orbit contains a representative in $K_{0}$, first note that $K_{i}=\phi^{i}(K_{0})$. Hence it suffices to prove that
$M=\bigcup_{i\in\mathbb{Z}}K_{i}.$
Viewing $M=\mathbb{R}\times S_{0}$ (as in Theorem 2.5), since the $S_{i}$ are Cauchy surfaces, each of them is the graph of a function $F_{i}\colon S_{0}\to\mathbb{R}$. We need to show that $F_{i}(x)\to\pm\infty$ when $i\to\pm\infty$, for all $x\in S_{0}$. Suppose that there is $x_{0}\in S_{0}$ such that $F_{i}(x_{0})\not\to\pm\infty$. Since $S_{i}\in I^{-}(S_{i+1})$, it follows that $F_{i}<F_{i+1}$. Thus, without loss of generality, $F_{i}(x_{0})\to C$ for $i\to+\infty$, where $C$ is a constant. Then, the points $x_{i}:=(F_{i}(x_{0}),x_{0})\in S_{i}$ converge to $x_{\infty}=(C,x_{0})$, and since the $S_{i}$ are totally timelike connected, $x_{i}\ll x_{i+1}$. We can construct a timelike curve $\gamma$ from $x_{0}$ to $x_{\infty}$ by concatenating maximizing geodesic segments $\gamma_{i}$ going from $x_{i}$ to $x_{i+1}$. Let $d\colon M\times M\to[0,\infty)$ be the Lorentzian distance (which is continuous by global hyperbolicity), and define
$D_{i}:=\inf\{d(p,q)\mid p\in S_{i},q\in S_{i+1}\}.$
By compactness of the $S_{i}$ and continuity of $d$, it follows that $D_{i}>0$. Moreover, since $d$ is invariant under isometries, $D_{i}$ is independent of $i$. Now, the Lorentzian length $L(\gamma)$ satisfies
$L(\gamma)=\sum_{i=0}^{\infty}L(\gamma_{i})>\sum_{i=0}^{\infty}D_{i}=\infty.$
It follows that $\gamma$ must be future inextendible, a contradiction.
It remains to show that $\tau$ is a Cauchy temporal function. We again distinguish between the cases of $G$ compact or non-compact. When $G$ is compact, it follows from (4) that $\tau$ has, up to an additive constant, the form
$\tau=\int_{G}\hat{\tau}\circ\phi\ d\mu(\phi).$
Let $\gamma\colon\mathbb{R}\to M$ be an inextendible causal curve (then, so is $\phi\circ\gamma$). Because $\hat{\tau}$ is Cauchy, we have that
$\hat{\tau}\circ\phi\circ\gamma(s)\to\pm\infty$
for every $\phi\in G$. By compactness of $G$, it then follows that $\tau(\gamma(s))\to\pm\infty$ as $s\to\pm\infty$, implying that $\tau$ is Cauchy. In the case that $G$ is non-compact, we use that the decomposition (6) implies that $(M,g)$ in fact satisfied the NOH. Indeed, any inextendible causal curve $\gamma$ must traverse all of the $S_{i}$’s, which are totally timelike connected, and it is easy to conclude that $I^{-}(\gamma)=M$. Then, it follows from Lemma 2.10 that $\tau$ must be Cauchy.
@@
Let $(M,g)$ be a causal spacetime satisfying the NFOH, and let $\tau\colon M\to\mathbb{R}$ be as in Theorem 3.2. Then
@@
[…]
@@ 3.2. Examples
Claim. For every $x\in\Sigma$, the vertical line $\mathbb{R}\times\{x\}\subset M$ is preserved by $\phi$.
@@
In the above proof, compactness of $\Sigma$ is essential. Indeed, Minkowski spacetime $\mathbb{R}^{1,n}$ is also a product spacetime, but its isometry group, the Poincaré group, includes boosts (which mix space and time), and hence is larger than $\mathbb{R}\times\mathsf{O}(n)$. Note also that product spacetimes with compact slices always satisfy the NOH. An example where the NOH is violated and the conclusion of Corollary 1.3 fails, is given by de Sitter spacetime $\mathsf{dS}^{4}$. On the one hand, $\mathsf{dS}^{4}$ can be represented as the warped product $\mathsf{dS}^{4}=\mathbb{R}^{1,0}\times_{\cosh^{2}}S^{3}$, implying that is is globally hyperbolic with compact Cauchy surfaces [4, pp. 183–184]. On the other hand, $\mathsf{dS}^{4}$ can also be represented as the pseudosphere in Minkowski spacetime $\mathbb{R}^{1,4}$. It follows that its isometry group is the Lorentz group $\mathsf{O}^{\uparrow}(1,4)$ [29, Prop. 9.8], which, again because it contains boosts, cannot be written in the form $\mathbb{R}\ltimes N$ with $N$ compact. While the obvious examples of spacetimes with isometry group $\mathbb{R}\ltimes N$ are the ones that admit a complete timelike Killing vector field (such as $\partial_{t}$ in the case of a product spacetime), the following example shows that they are not the only ones.
@@
Consider on $\mathbb{R}^{2}$ the Lorentzian metric $g:=f(y)dx^{2}-dy^{2}$ for some function $f\colon\mathbb{R}\to[0,1)$ with $f(y)=f(y+1)$. Let $M$ be the quotient of $\mathbb{R}^{2}$ by the map $(x,y)\mapsto(x+1,y+1)$. The metric $g$ descends to the quotient, which we can model as $\mathbb{R}\times[0,1]$ with the identifications $(0,y)\sim(1,y+1)$. Hence $M$ is homeomorphic to a cylinder. Because $f<1$, there are no closed causal curves, and in fact, $(M,g)$ is globally hyperbolic with $\{(s,s)\mid s\in[0,1]\}/\sim$ being a compact Cauchy surface. Moreover, $(M,g)$ satisfies the NOH. The coordinate vector field $\partial_{x}$ is complete, spacelike and Killing, with integral curves escaping to infinity. For generic $f$, it is the only Killing vector field; hence $\operatorname{Isom}^{\uparrow,0}(M,g)\cong\mathbb{R}$ even though $(M,g)$ does not admit a timelike Killing vector field.
@@
Finally, we argue that $\operatorname{Isom}^{\uparrow}(M,g)=L\ltimes N$, where $L$ is generated by $\phi$, is a semi-direct product, and is not isomorphic to $\mathbb{Z}\times N$. Clearly, $\phi$ does not commute with all elements in $N$. Concretely, if $\chi=(\chi_{1},\chi_{2})\in\mathsf{O}(2)\times\mathsf{O}(2)$, then
@@
Let $(M,g)$ be a causal spacetime satisfying the NFOH, and let $\tau\colon M\to\mathbb{R}$ be a Cauchy temporal function. Suppose $G$ is a group acting by isometries on $(M,g)$ such that, for every $t\in\mathbb{R}$ and one (equivalently, all) points $p\in\Sigma_{t}$, the orbit $G(p):=\{\phi(p)\mid\phi\in G\}$ equals $\Sigma_{t}$. Then
@@
Let $k\geq 2$ and let $\beta\colon\mathbb{R}\to\operatorname{GL}(k,\mathbb{R}),\ t\mapsto\beta_{t}$ be a one-parameter subgroup such that $\beta_{1}(\mathbb{Z}^{k})=\mathbb{Z}^{k}$ but $\beta_{1}\neq\operatorname{Id}$. On $\mathbb{R}\times\mathbb{R}^{k}$, consider the equivalence relation